Two angles are complementary. One angle is 15 degrees more than 2 times the other. Find the angles.
Answer by longjonsilver(2297)

complementary angles add to 90.

so let x=one angle
other angle is 15+2x

so x + 15+2x = 90
3x + 15 = 90
3x = 75
x = 75/3
x = 25 degrees

and other angle is therefore 65

jon.